OpenStudy (anonymous):

Calvin went cycling in San Francisco, which is extremely hilly. He can pedal up a hill at a speed of 12 mph, and down a hill at a speed of 36 mph. If he went up and down a hill (back to the same point), what is his average speed (in mph) for the entire journey?

OpenStudy (anonymous):

suppose the length from bottom of hill to the top is 'd'miles ; tht implies.. Time to go to top=d/12 Time to come from top=d/36 hence total tym taken = d/12+d/36=4d/36 hence AVG. speed= total dist / total tym =2d/(4d/36)=18 miles an hour